# Integrated modeling of boron powder injection for real-time plasma-facing component conditioning

### v3

@@ 3 Coupling impurity transport with erosion and deposition dynamics in WallDYN3D
In the WallDYN3D model, the evolution of surface density for materials on the reactor walls, crucial for assessing material behavior under plasma exposure, is described by differential equations that account for erosion, deposition, and material mixing processes. For this study, the WallDYN3D model was extended to include impurity sources located within the plasma domain resulting from the B powder particle evaporation. The B atomic sources provided by the EMC3-EIRENE/DIS calculations have been included in the WallDYN3D model as a charge state $qi$ resolved B influx of component $ei$ at each wall location $wk$: $\Gamma_{ei,wk,qi}^{Inj}$.
The total impurity influx into the plasma of element $ei$ from wall $wk$ is given by the sum of reflected, eroded/sublimated, and injected flux, which is summarized as such:
@@
where $m_{ei,wk,wj,qi}$ describes the charge state resolved redistribution and migration matrix, and $R_{ei,wk,qi}$, and $Y_{ei,em,wk,qm}$, describe the reflection and erosion yields, respectively. The governing equation for the areal density (atoms/m2) time evolution of wall surface element $\delta_{ei,wk}$ is defined as:
$\frac{d\delta_{ei,wk}}{dt}=\sum_{qi}\Gamma_{ei,wj,qi}^{In}-(\Gamma_{ei,wk}^{Ero}+\Gamma_{ei,wk}^{Refl})+\Gamma_{ei,wk}^{Bulk},$ (6)
@@ 4 Full-torus modeling of boron powder injection and transport with EMC3-EIRENE and DIS
A representative L-mode deuterium (D) plasma scenario for real-time wall conditioning in the DIII-D tokamak has been implemented using the EMC3-EIRENE code. This implementation utilizes a full-torus (360) grid based on the equilibrium #$\#179894$ at 3500 ms, specifically designed to fully capture the migration of boron powder particles and the resulting distributions of boron sources and fluxes. Previously, this scenario was used to model boron transport from a single atomic source [17]. In the current implementation, boron is introduced through powder injection, and the ablation and transport of this boron dust are simulated using the DIS code. A toroidal cross-section of the plasma grid domain for EMC3-EIRENE is shown in figure 1(a). The impurity powder source is situated at the tip of the upper small angle slot (SAS) divertor, where it is connected via a vertical tube to the powder dropper. The simulation accounts for a 5 cm diameter gap at this dropping location in the divertor geometry.
@@
The source rates were scaled to $10^{20}$ B/s and $10^{21}$ B/s to explore how experimental injection levels of a few milligrams per second up to 20 mg/s may affect the surface distribution and chemical composition (the plasma background was kept constant). Carbon, the predominant intrinsic impurity in the DIII-D tokamak due to its graphite PFCs, was included in the simulation. In the WallDYN3D simulations, the sputtering yields for C, both chemical and charge exchange, were set at 1%. The C flux from CX neutrals was estimated using redistribution matrices informed by EMC3-EIRENE background plasma parameters. Physical sputtering by the background D plasma was also considered, with yields calculated using the energy $3qT_{e}+2T_{i}$ based on the local plasma temperatures. The chemical and charge exchange sputtering for B were neglected. Multiple WallDYN3D time steps were carried out to model the evolution of the areal densities for B and C on the PFCs within the plasma domain. In each step, equations (5) and (6) are solved, resulting in effective recycling of the impurities.
@@
In WallDYN3D, impurity transport is treated kinetically by default [31]. A comparison between kinetic and fluid modeling was conducted to investigate potential differences in the long-range transport between the high and low-field sides of relatively light B particles. In this case, boron is sourced from the high field side target. The three-dimensional charge-state integrated boron redistribution matrices (see equation (5)) have been toroidally averaged, resulting in two-dimensional representations since this scenario assumed the wall geometry was toroidally symmetric. Figure 4(a) shows the wall elements and corresponding indices (IDX) in the poloidal plane. The boron redistribution matrices for both the kinetic (Figure 4b) and fluid (Figure 4c) approaches display the probability distribution of where impurities, once eroded, will travel and redeposit, specifically on the divertor wall elements. The diagonal terms indicate the dominating local redeposition of B on the inner ($IDX=6-12$) and outer divertor targets ($IDX=19-28$). The matrices also reveal an absence of transport from the inner to the outer target. However, values in the bottom right sections indicate additional transport from the outer to the inner target. On the inner target, less local redeposition occurs than on the outer target, with less eroded B (and C) returning to their origin (see diagonal elements in Figure 4b,c). This behavior is consistent in both the fluid and kinetic treatment qualitatively and quantitatively within the WallDYN3D model. The remaining differences between the kinetic and fluid models stem from how they handle impurity transport near the boron injection location. The fluid model assumes immediate equilibrium with the plasma, simplifying the process and leading to overestimated long-range transport. In contrast, the kinetic model accounts for the time needed for boron ions to accelerate and thermalize, resulting in more accurate, localized deposition.
@@ 6 Impact of boron source strength on mixed-material coating composition
The dynamics of the B-C surface composition are affected by erosion and reflection properties on wall elements. One-dimensional boron and carbon flux profiles at the toroidal DiMES location (dashed vertical line, [33]) are shown in figure 5 for time steps of 0, 1, 10, 20, and 400 seconds, with boron influx rates of $10^{20}$ and $10^{21}$ B/s (see poloidal wall index in figure 4a). The WallDYN3D calculations indicate that boron flux increases significantly due to reerosion, migration, and redeposition while carbon flux decreases. At the lower boron rate, boron flux to the high-field side target (inner peak) increases over time from $\approx 6\times 10^{19}$ to $1.2\times 10^{20}$ m-2s-1, and more dramatically on the outboard side (outer peak) from $\approx 3\times 10^{19}$ to $9\times 10^{20}$ m-2s-1. The carbon flux sees only a slight reduction. Conversely, at the higher boron rate ($10^{21}$ B/s), peak levels stabilize quickly, within 1 second at the inboard target, with carbon levels dropping by about 10%.
@@ Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards DE-AC02-09CH11466, DE-FC02-04ER54698, and DE-AC05-00OR22725. The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The data that support the findings of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.14176757.

### v4

@@ 3 Coupling impurity transport with erosion and deposition dynamics in WallDYN3D
In the WallDYN3D model, the evolution of surface density for materials on the reactor walls, crucial for assessing material behavior under plasma exposure, is described by differential equations that account for erosion, deposition, and material mixing processes. For this study, the WallDYN3D model was extended to include impurity sources located within the plasma domain resulting from the boron powder particle evaporation. The B atomic sources provided by the EMC3-EIRENE/DIS calculations are incorporated into the WallDYN3D model as a charge state ($qi$) resolved B injection influx ($\Gamma_{ei,wj,qi}^{Inj}$) at each wall segment ($wj$), where $wj$ represents a discretized location on the reactor wall.
The total impurity influx of component $ei$ on a wall element $wk$, defined as the sum of contributions from reflected, eroded/sublimated, and injected fluxes, is expressed as:
@@
where $m_{ei,wk,wj,qi}$ describes the charge state-resolved redistribution and migration matrix, and $R_{ei,wk,qi}$, and $Y_{ei,em,wk,qm}$, describe the reflection and erosion yields (including charge-exchange erosion), respectively. The governing equation for the areal density (atoms/m2) time evolution of wall surface element $\delta_{ei,wk}$ is defined as:
$\frac{d\delta_{ei,wk}}{dt}=\sum_{qi}\Gamma_{ei,wk,qi}^{In}-(\Gamma_{ei,wk}^{Ero}+\Gamma_{ei,wk}^{Refl})+\Gamma_{ei,wk}^{Bulk},$ (6)
@@ 4 Full-torus modeling of boron powder injection and transport with EMC3-EIRENE and DIS
A representative L-mode deuterium (D) plasma scenario for real-time wall conditioning in the DIII-D tokamak has been implemented using the EMC3-EIRENE code. This implementation utilizes a full-torus (360) grid based on the equilibrium #$\#179894$ at 3500 ms, specifically designed to fully capture the migration of boron powder particles and the resulting distributions of boron sources and fluxes. Previously, this scenario was used to model boron transport from a single atomic source [17]. In the current implementation, boron is introduced through powder injection, and the ablation and transport of this boron dust are simulated using the DIS code. A toroidal cross-section of the plasma grid domain for EMC3-EIRENE is shown in Figure 1(b). The impurity powder source is situated at the tip of the upper small angle slot (SAS) divertor, where it is connected via a vertical tube to the powder dropper. The simulation accounts for a 5 cm diameter gap at this dropping location in the divertor geometry.
@@
The source rates were scaled to $10^{20}$ B/s and $10^{21}$ B/s to explore how experimental injection levels of a few milligrams per second up to 20 mg/s may affect the surface distribution and chemical composition (the plasma background was kept constant). Carbon, the predominant intrinsic impurity in the DIII-D tokamak due to its graphite PFCs, was included in the simulation. In the WallDYN3D simulations, the sputtering yields for C, both chemical and charge exchange, were set at 1% [33]. CX neutral fluxes and average energies are estimated from EMC3-EIRENE background plasma parameters, assuming a Maxwellian-like energy distribution centered on the calculated average energy. Physical sputtering by the background D plasma was also considered, with yields calculated using the energy $3qT_{e}+2T_{i}$ based on the local plasma temperatures. The chemical and charge exchange sputtering for B were neglected. Multiple WallDYN3D time steps were carried out to model the evolution of the areal densities for B and C on the PFCs within the plasma domain. In each step, equations (5) and (6) are solved, resulting in effective recycling of the impurities.
@@
In WallDYN3D, impurity transport is treated kinetically by default [31]. A comparison between kinetic and fluid modeling was conducted to investigate potential differences in the long-range transport between the high and low-field sides of relatively light B particles. In this case, boron is sourced from the high field side target. The three-dimensional charge-state integrated boron redistribution matrices (see equation (5)) have been toroidally averaged, resulting in two-dimensional representations since this scenario assumed the wall geometry was toroidally symmetric. Figure 4(a) shows the wall elements and corresponding indices (IDX) in the poloidal plane. The boron redistribution matrices for both the kinetic (Figure 4(b)) and fluid (Figure 4(c)) approaches display the probability distribution of where impurities, once eroded, will travel and redeposit, specifically on the divertor wall elements. The diagonal terms indicate the dominating local redeposition of B on the inner ($IDX=6-12$) and outer divertor targets ($IDX=19-28$). The matrices also reveal an absence of transport from the inner to the outer target. However, values in the bottom right sections indicate additional transport from the outer to the inner target. On the inner target, less local redeposition occurs than on the outer target, with less eroded B (and C) returning to their origin (see diagonal elements in Figure 4(b,c)). This behavior is consistent in both the fluid and kinetic treatment qualitatively and quantitatively within the WallDYN3D model. The remaining differences between the kinetic and fluid models stem from how they handle impurity transport near the boron injection location. The fluid model assumes immediate equilibrium with the plasma, simplifying the process and leading to overestimated long-range transport. In contrast, the kinetic model accounts for the time needed for boron ions to accelerate and thermalize, resulting in more accurate, localized deposition.
@@ 6 Impact of boron source strength on mixed-material coating composition
The dynamics of the B-C surface composition are affected by erosion and reflection properties on wall elements. One-dimensional boron and carbon flux profiles at the toroidal DiMES location (dashed vertical line, [34]) are shown in Figure 5 for time steps of 0, 1, 10, 20, and 400 seconds, with boron influx rates of $10^{20}$ and $10^{21}$ B/s (see poloidal wall index in Figure 4(a)). The WallDYN3D calculations indicate that boron flux increases significantly due to reerosion, migration, and redeposition while carbon flux decreases. At the lower boron rate, boron flux to the high-field side target (inner peak) increases over time from $\approx 6\times 10^{19}$ to $1.2\times 10^{20}$ m-2s-1, and more dramatically on the outboard side (outer peak) from $\approx 3\times 10^{19}$ to $9\times 10^{20}$ m-2s-1. The carbon flux sees only a slight reduction. Conversely, at the higher boron rate ($10^{21}$ B/s), peak levels stabilize quickly, within 1 second at the inboard target, with carbon levels dropping by about 10%.
@@ Acknowledgements
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards DE-AC02-09CH11466, DE-FC02-04ER54698, and DE-AC05-00OR22725. The United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The data that support the findings of this study are openly available in Zenodo at https://zenodo.org/records/14176757.